A portable electric water heater transfers 255 watts (W) of power to 5.5 L of water, where 1 W=1J/s. A portable electric water heater transfers 255 watts (W) of power to 5.5 L of water, where 1 W=1J/s.
1
Expert's answer
2019-10-21T07:16:15-0400

"T_1=25\degree C;T_2=42\degree C;m=1000\times 5.5 \times 10^{-3}=" "5.5 \ kg"

"C" for water"=4.184 \times 10^3 J/Kg"

Energy required to raise this temperature"=" "mC(T_2-T_1)=5.5\times 4.184 \times 10^3\times (42-25)=" "391.204\times 10^3 J"

Power given"=255 J\/sec"

Time required"=\frac{Energy}{Power}=\frac{391.204\times 10^3}{255}=" "154.133 \ sec=\frac{154.133}{60}=2.585 \ min"